6. Which is relatively better: a score of 76 on a psychology test or a score of 33 on an economics test?
Scores on the psychology test have a mean of 88 and a standard deviation of 10. Scores on the economics test have a mean of 47 and a standard deviation of 5
Choose the correct answer below.
a. The economics test score is relatively better because its z score is less than the z score for the psychology test score.
b. The psychology test score is relatively better because its z score is greater than the z score for the economics test score.
c. The economics test score is relatively better because its z score is greater than the z score of the psychology test.
d. The psychology test score is relatively better because its z score is less than the z score for the economics test score.

Answer by CubeyThePenguin(3113)

z = (x - mu)/(sigma)

z = z score
x = observed value
mu = mean
sigma = standard deviation

psychology:
z = (76 - 88)/(10) = -1.2

economics:
z = (33 - 47)/5 = -2.8

The psychology test score is relatively better because its z-score is greater than the z-score of the economics test.